# LanPaint: Training-Free Diffusion Inpainting with Asymptotically Exact and Fast Conditional Sampling

### v2

@@ Appendix B The Denoising Diffusion Probabilistic Model (DDPMs)
The backward diffusion process is the conjugate of the forward process. While the forward process evolves $p_{t}$ toward $\mathcal{N}(\mathbf{0},I)$, the backward process reverses this evolution, restoring $\mathcal{N}(\mathbf{0},I)$ to $p_{t}$. To derive it, we know from previous section that Langevin dynamics equation 20 acts as an "identity" operation on a distribution. Thus, the composition of forward and backward processes, at time $t$, must yield the Langevin dynamics for $p_{t}(\mathbf{x})$.
@@
where $\mathbf{s}(\mathbf{x},t)=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is the score function of $p_{t}(\mathbf{x})$. The "Forward" part corresponds to the forward diffusion process equation 22, effectively increasing the forward diffusion time $t$ by $d\tau$, bringing the distribution to $p_{t+d\tau}(\mathbf{x})$. Since the forward and backward components combine to form an "identity" operation, the "Backward" part in equation 23 must reverse the forward process—decreasing the forward diffusion time $t$ by $d\tau$ and restoring the distribution back to $p_{t}(\mathbf{x})$.
Now we can define the backward process according to the backward part in equation 23, and a backward diffusion time $t^{\prime}$ different from the forward diffusion time $t$:
@@
In practice, the forward OU process equation 22 is numerically discretized into the variance-preserving (VP) form (Song et al., 2020b):
@@
where $i=1,\cdots,n$ is the number of the time step, $\beta_{i}$ is the step size of each time step, $\mathbf{x}_{i}$ is image at $i$th time step with time $t_{i}=\sum_{j=0}^{i-1}\beta_{j}$, $\bm{\epsilon}_{i}$ is standard Gaussian random variable. The time step size usually takes the form $\beta_{i}=\frac{i(b_{2}-b_{1})}{n-1}+b_{1}$ where $b_{1}=10^{-4}$ and $b_{2}=0.02$. Note that our interpretation of $\beta$ differs from that in Song et al. (2020b), treating $\beta$ as a varying time-step size to solve the autonomous SDE equation 22 instead of a time-dependent SDE. Our interpretation holds as long as every $\beta_{i}^{2}$ is negligible and greatly simplifies future analysis. The discretized OU process equation 30 adds a small amount of Gaussian noise to the image at each time step $i$, gradually contaminating the image until $\mathbf{x}_{n}\sim\mathcal{N}(\mathbf{0},I)$.
Training a DDPM aims to recover the original image $x_{0}$ from one of its contaminated versions $x_{i}$. In this case equation 30 could be rewritten into the forward diffusion process
@@
[…]
The backward diffusion process is used to sample from the DDPM by removing the noise of an image step by step. It is the time reversed version of the OU process, starting at $x_{0^{\prime}}\sim\mathcal{N}(\mathbf{x}|\mathbf{0},I)$, using the reverse of the OU process equation 26. In practice, the backward diffusion process is discretized into
@@
where $i^{\prime}=0,\cdots,n$ is the number of the backward time step, $\mathbf{x}_{i^{\prime}}$ is image at $i^{\prime}$th backward time step with time $t_{i^{\prime}}^{\prime}=\sum_{j=0}^{i^{\prime}-1}\beta_{n-1-j}=T-t_{n-i^{\prime}}$. This discretization is consistent with equation 26 as long as $\beta_{i}^{2}$ are negligible. The score function $\mathbf{s}(\mathbf{x}_{i^{\prime}},T-t^{\prime}_{i^{\prime}})$ is generally modeled by a neural network and trained with a denoising objective.
@@
DDPM is trained to removes the noise $\bar{\bm{\epsilon}}_{i}$ from $\mathbf{x}_{i}$ in equation 31, by training a denoising neural network $\bm{\epsilon}_{\theta}(\mathbf{x},t_{i})$ to predict and remove the noise $\bar{\bm{\epsilon}}_{i}$. This means that DDPM minimizes the denoising objective (Ho et al., 2020):
@@
Now we show that $\bm{\epsilon}_{\theta}$ trained with the above objective is proportional to the score function $\mathbf{s}$. Note that the equation 31 tells us that the distribution of $\mathbf{x}_{i}$ given $\mathbf{x}_{0}$ is a Gaussian distribution
@@
and the noise $\bar{\bm{\epsilon}}_{i}$ in equation 31 is directly proportional to the score function
@@
The equation 36 is an important property. It tells us that the noise $\bar{\bm{\epsilon}}_{i}$ is directly related to a conditional score function. This conditional score function is connected to the score function $\mathbf{s}(\mathbf{x},t)$ through the following equation:
@@
Substituting equation 36 into equation 34 and utilizing equation 37, we could derive that equation 34 is equivalent to a denoising score matching objective
@@
Therefore the denoising neural network is actually a scaled estimate of the score function $\mathbf{s}(\mathbf{x},t)$, hence could be inserted into the backward sampling process equation 33 to generate images.
Appendix C The ODE Based Backward Diffusion Process
The backward diffusion process equation 24 is not the only reverse process for the forward process equation 22. We can derive a deterministic ordinary differential equation (ODE) as an alternative, removing the stochastic term $d\mathbf{W}$ in the reverse process.
To obtain this ODE reverse process, consider the Langevin dynamics equation 23 with a rescaled time ($d\tau\rightarrow\frac{1}{2}d\tau$):
$\begin{split}d\mathbf{x}_{\tau}&=\frac{1}{2}\mathbf{s}(\mathbf{x}_{\tau},t)d\tau+\,d\mathbf{W}_{\tau},\\ &=\underbrace{-\frac{1}{2}\mathbf{x}_{\tau}d\tau+d\mathbf{W}_{\tau}}_{\text{Forward}}+\underbrace{\frac{1}{2}\mathbf{x}_{t}d\tau+\frac{1}{2}\mathbf{s}(\mathbf{x}_{\tau},t)d\tau}_{\text{Backward}},\end{split}$ (40)
Following the same logic used to derive the backward diffusion process equation 26, we extract from this splitting the backward ODE (known as the probability flow ODE Song et al. (2020b)):
@@ Appendix C The ODE Based Backward Diffusion Process
where $t^{\prime}\in[0,T]$ is backward time, and $\mathbf{s}(\mathbf{x},t)=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x})$ is the score function of the density of $\mathbf{x}_{t}$ in the forward process. This ODE maintains the same forward-backward duality as the SDE reverse process equation 26.
@@
Appendix D Three Notations of Diffusion Models
@@
To simplify notation, we now use continuous time $t$ and its corresponding state $\mathbf{x}_{t}$ (as in equation 22), rather than discrete notations like $t_{i}$ and $\mathbf{x}_{i}$.
@@
The DDPMs introduced in the previous section are called ’variance-preserving’ models. This name originates from the forward process equation 31: if the clean images $\mathbf{x}_{0}$ are normalized such that $\mathrm{Cov}(\mathbf{x}_{0},\mathbf{x}_{0})=I$, then this covariance is preserved at any time $t_{i}$, with $\mathrm{Cov}(\mathbf{x}_{i},\mathbf{x}_{i})=I$.
The forward diffusion process equation 31 in continuous time $t$ is:
@@
where $\bar{\alpha}_{t}=e^{-t}$ (from equation 32) and $\bar{\bm{\epsilon}}_{t}\sim\mathcal{N}(0,\mathbf{I})$.
@@
    where $t^{\prime}\in[0,T]$ is reversed time, and the score function $\mathbf{s}(\mathbf{x},t)=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is learned via the denoising objective equation 34 and equation 39.
While we previously trained the denoising network $\bm{\epsilon}_{\theta}$ using objective equation 34 (related to the score function via equation 39), we can alternatively model the score function $\mathbf{s}_{\theta}$ directly. By substituting $\bm{\epsilon}_{\theta}$ with $\mathbf{s}_{\theta}$ and dropping the $\sqrt{1-\bar{\alpha}_{t}}$ scaling factor, we obtain the equivalent score-based objective:
@@
where $\mathbf{x}_{t}$ follows equation 42. This represents an equivalent but reweighted version of the original denoising objective equation 34.
@@
Rewriting the VP forward equation 42 in VE notation yields:
@@
To directly model $\bm{\epsilon}_{\theta}(\mathbf{z},\sigma)$, we adapt the denoising objective equation 34 to VE coordinates by replacing $\mathbf{x}_{t}$ with $\mathbf{z}_{\sigma}$:
@@
where $\sigma_{max}=\sqrt{(1-\bar{\alpha}_{T})/\bar{\alpha}_{T}}$ and $\mathbf{z}_{\sigma}$ follows equation 47. This preserves the denoising objective’s structure while operating in VE space.
@@
While often presented as a distinct framework from DDPMs, rectified flows are mathematically equivalent Gao et al. (2025) to DDPMs. We now provide a much simpler proof via the transformations:
@@
Rewriting the VE forward process equation 47 in RF coordinates yields:
@@
To directly model $\mathbf{v}_{\theta}(\mathbf{r}_{s^{\prime}},s^{\prime})$, we transform the denoising objective equation 50 by substituting $\bm{\epsilon}_{\theta}$ with $\mathbf{v}_{\theta}$ and inserting the RF forward process equation 52 into equation 50, while removing a constant scaling factor $(1-s)$. This yields the flow matching objective:
@@
where $\mathbf{r}_{s}$ follows equation 52. This represents a re-weighted equivalent of the denoising objective equation 34, interpreted in the flow matching framework where $\bar{\bm{\epsilon}}$ corresponds to the endpoint $\mathbf{r}_{1}$ and $\mathbf{v}_{\theta}$ models the velocity field transporting $\mathbf{r}_{0}$ to $\mathbf{r}_{1}$.
In summary, the three notations (VP, VE, and RF) are mutually transformable through the mappings defined in equation 46 and equation 51. This equivalence enables a practical LanPaint implementation strategy: we can design LanPaint using any single notation (such as VP) and automatically extend it to other frameworks by applying these transformations.
Appendix E Stationary Distribution of Langevin Dynamics with the BIG score
In this appendix, we prove that the BiG score Langevin Dynamics defined in equation 13 converges to the target distribution
$\pi_{t}(\mathbf{x},\mathbf{y})\propto p_{t}(\mathbf{x}\mid\mathbf{y})\,\frac{p_{t}(\mathbf{y}\mid\mathbf{y}_{o})^{1+\lambda}}{p_{t}(\mathbf{y})^{\lambda}}+o(\sqrt{1-\bar{\alpha}_{t}}).$
@@ Appendix E Stationary Distribution of Langevin Dynamics with the BIG score
$\mathbf{g}_{\lambda}(\mathbf{x},\mathbf{y},t)=-\left((1+\lambda)\,\frac{\mathbf{y}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}+\lambda\,\mathbf{s}_{\mathbf{y}}(\mathbf{x},\mathbf{y},t)\right).$
E.1 Idealized SDE for the Target Distribution
@@
$\pi^{*}(\mathbf{x}_{t},\mathbf{y}_{t})\propto p(\mathbf{x}_{t}\mid\mathbf{y}_{t})\,\frac{p(\mathbf{y}_{t}\mid\mathbf{y}_{o})^{1+\lambda}}{p(\mathbf{y}_{t})^{\lambda}}.$
@@
$\mathbf{s}^{*}(\mathbf{x},\mathbf{y},t)=\nabla_{\mathbf{x},\mathbf{y}}\left[\log p_{t}(\mathbf{x}\mid\mathbf{y})+(1+\lambda)\log p_{t}(\mathbf{y}\mid\mathbf{y}_{o})-\lambda\log p_{t}(\mathbf{y})\right].$
@@
$\mathbf{s}^{*}_{\mathbf{x}}(\mathbf{x},\mathbf{y},t)$ $=\mathbf{s}_{\mathbf{x}}(\mathbf{x},\mathbf{y},t),$
@@
$\mathbf{s}^{*}_{\mathbf{y}}=(1+\lambda)\,\nabla_{\mathbf{y}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{y}_{t})-(1+\lambda)\,\frac{\mathbf{y}_{t}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}-\lambda\,\nabla_{\mathbf{y}_{t}}\log p(\mathbf{x}_{t},\mathbf{y}_{t}).$
E.2 Comparison with BiG score Drift Term

### v3

@@ Appendix C The Denoising Diffusion Probabilistic Model (DDPMs)
The backward diffusion process is the conjugate of the forward process. While the forward process evolves $p_{t}$ toward $\mathcal{N}(\mathbf{0},I)$, the backward process reverses this evolution, restoring $\mathcal{N}(\mathbf{0},I)$ to $p_{t}$. To derive it, we know from previous section that Langevin dynamics Eq.22 acts as an "identity" operation on a distribution. Thus, the composition of forward and backward processes, at time $t$, must yield the Langevin dynamics for $p_{t}(\mathbf{x})$.
@@
where $\mathbf{s}(\mathbf{x},t)=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is the score function of $p_{t}(\mathbf{x})$. The "Forward" part corresponds to the forward diffusion process Eq.24, effectively increasing the forward diffusion time $t$ by $d\tau$, bringing the distribution to $p_{t+d\tau}(\mathbf{x})$. Since the forward and backward components combine to form an "identity" operation, the "Backward" part in Eq.25 must reverse the forward process—decreasing the forward diffusion time $t$ by $d\tau$ and restoring the distribution back to $p_{t}(\mathbf{x})$.
Now we can define the backward process according to the backward part in Eq.25, and a backward diffusion time $t^{\prime}$ different from the forward diffusion time $t$:
@@
In practice, the forward OU process Eq.24 is numerically discretized into the variance-preserving (VP) form (Song et al., 2020c):
@@
where $i=1,\cdots,n$ is the number of the time step, $\beta_{i}$ is the step size of each time step, $\mathbf{x}_{i}$ is image at $i$th time step with time $t_{i}=\sum_{j=0}^{i-1}\beta_{j}$, $\bm{\epsilon}_{i}$ is standard Gaussian random variable. The time step size usually takes the form $\beta_{i}=\frac{i(b_{2}-b_{1})}{n-1}+b_{1}$ where $b_{1}=10^{-4}$ and $b_{2}=0.02$. Note that our interpretation of $\beta$ differs from that in (Song et al., 2020c), treating $\beta$ as a varying time-step size to solve the autonomous SDE Eq.24 instead of a time-dependent SDE. Our interpretation holds as long as every $\beta_{i}^{2}$ is negligible and greatly simplifies future analysis. The discretized OU process Eq.32 adds a small amount of Gaussian noise to the image at each time step $i$, gradually contaminating the image until $\mathbf{x}_{n}\sim\mathcal{N}(\mathbf{0},I)$.
Training a DDPM aims to recover the original image $x_{0}$ from one of its contaminated versions $x_{i}$. In this case Eq.32 could be rewritten into the forward diffusion process
@@
$\lim_{\max_{j}\beta_{j}\xrightarrow[]{}0}\bar{\alpha}_{i}\xrightarrow[]{}e^{-t_{i}}.$ (34)
The backward diffusion process is used to sample from the DDPM by removing the noise of an image step by step. It is the time reversed version of the OU process, starting at $x_{0^{\prime}}\sim\mathcal{N}(\mathbf{x}|\mathbf{0},I)$, using the reverse of the OU process Eq.28. In practice, the backward diffusion process is discretized into
@@
where $i^{\prime}=0,\cdots,n$ is the number of the backward time step, $\mathbf{x}_{i^{\prime}}$ is image at $i^{\prime}$th backward time step with time $t_{i^{\prime}}^{\prime}=\sum_{j=0}^{i^{\prime}-1}\beta_{n-1-j}=T-t_{n-i^{\prime}}$. This discretization is consistent with Eq.28 as long as $\beta_{i}^{2}$ are negligible. The score function $\mathbf{s}(\mathbf{x}_{i^{\prime}},T-t^{\prime}_{i^{\prime}})$ is generally modeled by a neural network and trained with a denoising objective.
@@
DDPM is trained to removes the noise $\bar{\bm{\epsilon}}_{i}$ from $\mathbf{x}_{i}$ in Eq.33, by training a denoising neural network $\bm{\epsilon}_{\theta}(\mathbf{x},t_{i})$ to predict and remove the noise $\bar{\bm{\epsilon}}_{i}$. This means that DDPM minimizes the denoising objective (Ho et al., 2020):
@@
Now we show that $\bm{\epsilon}_{\theta}$ trained with the above objective is proportional to the score function $\mathbf{s}$. Note that the Eq.33 tells us that the distribution of $\mathbf{x}_{i}$ given $\mathbf{x}_{0}$ is a Gaussian distribution
@@
and the noise $\bar{\bm{\epsilon}}_{i}$ in Eq.33 is directly proportional to the score function
@@
The Eq.38 is an important property. It tells us that the noise $\bar{\bm{\epsilon}}_{i}$ is directly related to a conditional score function. This conditional score function is connected to the score function $\mathbf{s}(\mathbf{x},t)$ through the following equation:
@@
Substituting Eq.38 into Eq.36 and utilizing Eq.39, we could derive that Eq.36 is equivalent to a denoising score matching objective
@@
Therefore the denoising neural network is actually a scaled estimate of the score function $\mathbf{s}(\mathbf{x},t)$, hence could be inserted into the backward sampling process Eq.35 to generate images.
Appendix D The ODE Based Backward Diffusion Process
The backward diffusion process Eq.26 is not the only reverse process for the forward process Eq.24. We can derive a deterministic ordinary differential equation (ODE) as an alternative, removing the stochastic term $d\mathbf{W}$ in the reverse process.
To obtain this ODE reverse process, consider the Langevin dynamics Eq.25 with a rescaled time ($d\tau\rightarrow\frac{1}{2}d\tau$):
$\begin{split}d\mathbf{x}_{\tau}&=\frac{1}{2}\mathbf{s}(\mathbf{x}_{\tau},t)d\tau+\,d\mathbf{W}_{\tau},\\ &=\underbrace{-\frac{1}{2}\mathbf{x}_{\tau}d\tau+d\mathbf{W}_{\tau}}_{\text{Forward}}+\underbrace{\frac{1}{2}\mathbf{x}_{\tau}d\tau+\frac{1}{2}\mathbf{s}(\mathbf{x}_{\tau},t)d\tau}_{\text{Backward}},\end{split}$ (42)
Following the same logic used to derive the backward diffusion process Eq.28, we extract from this splitting the backward ODE (known as the probability flow ODE (Song et al., 2020c)):
@@ Appendix D The ODE Based Backward Diffusion Process
where $t^{\prime}\in[0,T]$ is backward time, and $\mathbf{s}(\mathbf{x},t)=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x})$ is the score function of the density of $\mathbf{x}_{t}$ in the forward process. This ODE maintains the same forward-backward duality as the SDE reverse process Eq.28.
@@
Appendix E Three Notations of Diffusion Models
@@
To simplify notation, we now use continuous time $t$ and its corresponding state $\mathbf{x}_{t}$ (as in Eq.24), rather than discrete notations like $t_{i}$ and $\mathbf{x}_{i}$.
@@
The DDPMs introduced in the previous section are called ’variance-preserving’ models. This name originates from the forward process Eq.33: if the clean images $\mathbf{x}_{0}$ are normalized such that $\mathrm{Cov}(\mathbf{x}_{0},\mathbf{x}_{0})=I$, then this covariance is preserved at any time $t_{i}$, with $\mathrm{Cov}(\mathbf{x}_{i},\mathbf{x}_{i})=I$.
The forward diffusion process Eq.33 in continuous time $t$ is:
@@
where $\bar{\alpha}_{t}=e^{-t}$ (from Eq.34) and $\bar{\bm{\epsilon}}_{t}\sim\mathcal{N}(0,\mathbf{I})$.
@@
    where $t^{\prime}\in[0,T]$ is reversed time, and the score function $\mathbf{s}(\mathbf{x},t)=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is learned via the denoising objective Eq.36 and Eq.41.
While we previously trained the denoising network $\bm{\epsilon}_{\theta}$ using objective Eq.36 (related to the score function via Eq.41), we can alternatively model the score function $\mathbf{s}_{\theta}$ directly. By substituting $\bm{\epsilon}_{\theta}$ with $\mathbf{s}_{\theta}$ and dropping the $\sqrt{1-\bar{\alpha}_{t}}$ scaling factor, we obtain the equivalent score-based objective:
@@
where $\mathbf{x}_{t}$ follows Eq.44. This represents an equivalent but reweighted version of the original denoising objective Eq.36.
@@
Rewriting the VP forward Eq.44 in VE notation yields:
@@
To directly model $\bm{\epsilon}_{\theta}(\mathbf{z},\sigma)$, we adapt the denoising objective Eq.36 to VE coordinates by replacing $\mathbf{x}_{t}$ with $\mathbf{z}_{\sigma}$:
@@
where $\sigma_{max}=\sqrt{(1-\bar{\alpha}_{T})/\bar{\alpha}_{T}}$ and $\mathbf{z}_{\sigma}$ follows Eq.49. This preserves the denoising objective’s structure while operating in VE space.
@@
While often presented as a distinct framework from DDPMs, rectified flows are mathematically equivalent (Gao et al., 2025) to DDPMs. We now provide a much simpler proof via the transformations:
@@
Rewriting the VE forward process Eq.49 in RF coordinates yields:
@@
To directly model $\mathbf{v}_{\theta}(\mathbf{r}_{s^{\prime}},s^{\prime})$, we transform the denoising objective Eq.52 by substituting $\bm{\epsilon}_{\theta}$ with $\mathbf{v}_{\theta}$ and inserting the RF forward process Eq.54 into Eq.52, while removing a constant scaling factor (1s)(1-s). This yields the flow matching objective:
@@
where $\mathbf{r}_{s}$ follows Eq.54. This represents a re-weighted equivalent of the denoising objective Eq.36, interpreted in the flow matching framework where $\bar{\bm{\epsilon}}$ corresponds to the endpoint $\mathbf{r}_{1}$ and $\mathbf{v}_{\theta}$ models the velocity field transporting $\mathbf{r}_{0}$ to $\mathbf{r}_{1}$.
In summary, the three notations (VP, VE, and RF) are mutually transformable through the mappings defined in Eq.48 and Eq.53. This equivalence enables a practical LanPaint implementation strategy: we can design LanPaint using any single notation (such as VP) and automatically extend it to other frameworks by applying these transformations.
Appendix F Stationary Distribution of Langevin Dynamics with the BiG score
In this appendix, we prove that the BiG score Langevin Dynamics defined in Eq.13 converges to the target distribution
$\pi_{t}(\mathbf{x},\mathbf{y})\propto p_{t}(\mathbf{x}\mid\mathbf{y})\,\frac{p_{t}(\mathbf{y}\mid\mathbf{y}_{o})^{1+\lambda}}{p_{t}(\mathbf{y})^{\lambda}}+o(\sqrt{1-\bar{\alpha}_{t}}).$ (58)
@@ Appendix F Stationary Distribution of Langevin Dynamics with the BiG score
$\mathbf{g}_{\lambda}(\mathbf{x},\mathbf{y},t)=-\left((1+\lambda)\,\frac{\mathbf{y}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}+\lambda\,\mathbf{s}_{\mathbf{y}}(\mathbf{x},\mathbf{y},t)\right).$ (60)
F.1 Idealized SDE for the Target Distribution
@@
$\pi^{*}(\mathbf{x}_{t},\mathbf{y}_{t})\propto p(\mathbf{x}_{t}\mid\mathbf{y}_{t})\,\frac{p(\mathbf{y}_{t}\mid\mathbf{y}_{o})^{1+\lambda}}{p_{t}(\mathbf{y})^{\lambda}}.$ (61)
@@
$\mathbf{s}^{*}(\mathbf{x},\mathbf{y},t)=\nabla_{\mathbf{x},\mathbf{y}}\left[\log p_{t}(\mathbf{x}\mid\mathbf{y})+(1+\lambda)\log p_{t}(\mathbf{y}\mid\mathbf{y}_{o})-\lambda\log p_{t}(\mathbf{y})\right].$ (62)
@@
$\mathbf{s}^{*}_{\mathbf{x}}(\mathbf{x},\mathbf{y},t)$ $=\mathbf{s}_{\mathbf{x}}(\mathbf{x},\mathbf{y},t),$ (64)
@@
$\mathbf{s}^{*}_{\mathbf{y}}=(1+\lambda)\,\nabla_{\mathbf{y}_{t}}\log p(\mathbf{x}_{t}\mid\mathbf{y}_{t})-(1+\lambda)\,\frac{\mathbf{y}_{t}-\sqrt{\bar{\alpha}_{t}}\,\mathbf{y}_{o}}{1-\bar{\alpha}_{t}}-\lambda\,\nabla_{\mathbf{y}_{t}}\log p(\mathbf{x}_{t},\mathbf{y}_{t}).$ (65)
F.2 Comparison with BiG score Drift Term